a rectangle having a perimeter of 182 yards has a width 20 yards shorter than twice its length what is the area of this rectangle??
Answer by rfer(16322)

2x+2(2x-20)=182
2x+4x-40=182
6x=222
x=37
2x-20=54
-----------
A=37*54
A=1998 sq yds